# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v1

@@ Abstract
The convergence of two technological trajectories, large language models that automate cognitive labor and deep reinforcement learning agents that automate physical labor, implies the near-complete elimination of human employment as historically understood. The universal approximation theorem establishes that any continuous function can be represented by a sufficiently expressive neural network; foundational results in deep reinforcement learning establish that optimal policies exist for sequential decision tasks. Together, these results imply that all labor, white-collar and blue-collar, is in principle automatable, though significant engineering challenges remain. The critical question is not whether full automation will arrive, but who will control it. This paper introduces anarcho-accelerationism: a sociotechnical framework for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care rather than corporate extraction. Synthesizing anarchist political theory, commons-based peer production, and critical technology studies, we propose the Liberation Stack, a layered architecture of energy, manufacturing, food, communication, knowledge, and governance commons built on existing open-source technologies. We demonstrate that anarcho-accelerationism functions as a bridge-building framework compatible with liberalism, socialism, environmentalism, feminism, cooperativism, and the hacker ethic. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, and guifi.net shows that commons-based systems already operate at scale. We further argue that full automation renders money obsolete, making Universal Basic Income an insufficient response; we propose instead Universal Desired Resources (UDR), a post-monetary design principle where every person requests what they need from the robotic commons, constrained only by ecological sustainability. Drawing on prior work demonstrating the independence of phenomenal consciousness from computational intelligence, we establish that delegating all labor to non-conscious machines is not exploitation but care at civilizational scale, and that moral policy for the commons can be studied and refined through deep reinforcement learning. We conclude that one of the most important engineering challenges of the 21st century is building the decentralized infrastructure through which robotic automation serves all of humanity, enabling a society centered on care, love, and the common good.
@@
Keywords: commons-based peer production; federated systems; cooperative platforms; sociotechnical design; mutual aid; digital sovereignty; open-source AI; mesh networks; platform cooperativism; democratic technology; universal approximation; deep reinforcement learning; full automation; universal desired resources; machine consciousness; computational ethics
@@ 1 Introduction: The Convergence Window
The central claim of this paper is that anarcho-accelerationism is not a partisan ideology but a bridge-building framework. It finds genuine common ground with every major political and intellectual tradition of our time (Section˜3), while offering a technical architecture (Section˜4) grounded in empirical evidence (Section˜5). We believe this matters for computer science because we are the ones building the infrastructure. Every protocol we design, every platform we deploy, every algorithm we optimize embodies a political choice about who has power and who does not (Winner, 1980; Lessig, 2006). We can choose differently.
@@ 2 The Automation Imperative: Why This Framework, Why Now
The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The Liberation Stack (Section˜4) describes existing technologies, but the trajectory points further. Advances in additive manufacturing (3D printing at molecular scales), synthetic biology (DNA-programmable organisms and materials), and recursive AI self-improvement (where AI systems design better AI systems, better robots, better energy systems, and better versions of themselves) converge toward a productive infrastructure capable of synthesizing any physical object, any material, any energy configuration, bounded only by the laws of physics rather than the limits of human labor or capital. 3D bioprinters already fabricate living tissue; DNA origami techniques fold genetic material into nanoscale structures; large language models already write the code that trains the next generation of models. The logical endpoint is not merely automation of existing jobs but the creation of a universal fabrication commons: a decentralized network of robotic systems capable of producing, on demand, whatever any community or individual needs, from food to medicine to housing to energy infrastructure.
@@
Taken together, these four consequences yield a vision that goes well beyond conventional proposals for technology-assisted social reform. We are not proposing a better welfare state, a more humane capitalism, or a fairer distribution of existing wealth. We are proposing that the combination of decentralized full automation, commons governance, and the ontological impossibility of machine suffering makes possible, for the first time in history, a society in which no conscious being is compelled to perform instrumental labor, and every conscious being has access to whatever they need to live a life of care, creativity, and meaning. The remainder of this paper provides the political bridges (Section˜3), the technical architecture (Section˜4), and the empirical evidence (Section˜5) for building this society.
@@ 3 The Bridge-Building Thesis
Faith traditions Common good Shared dignity, mutual aid Operationalization through commons governance
@@ 4 The Liberation Stack: A Technical Architecture
Anarcho-accelerationism is not merely a philosophical position. It proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜4. Each layer is composed of existing, mature, open-source technologies.
@@
The Liberation Stack is designed for compositional sovereignty: each layer can be adopted independently, but the layers reinforce each other. A community with energy sovereignty has more bargaining power. A community with communication sovereignty cannot be silenced. A community with governance tools can make collective decisions about all other layers. Together, the layers form a complete alternative infrastructure that does not depend on either state provision or market availability. Table˜4 summarizes the architecture.

### v2

@@ Abstract
The convergence of large language models that automate cognitive labor and deep reinforcement learning agents that automate physical labor implies the near-complete elimination of human employment as historically understood. The universal approximation theorem and foundational results in deep reinforcement learning establish that all labor is in principle automatable. The critical question is not whether full automation will arrive, but who will control it. This paper introduces anarcho-accelerationism: a sociotechnical framework for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. We propose the Liberation Stack, a layered architecture of energy, manufacturing, food, communication, knowledge, and governance commons, powered by frontier clean energy technologies (perovskite tandem solar, enhanced geothermal, nuclear fusion, iron-air batteries) within an accelerationist ecologism that achieves sustainability through technological abundance rather than degrowth. We demonstrate that anarcho-accelerationism builds bridges to liberalism, socialism, environmentalism, feminism, cooperativism, faith traditions, and the hacker ethic, and that it respects freedom of faith on principled grounds: the ontological scope of theology escapes the epistemological reach of artificial intelligence, as argued through the frameworks of Saint Teresa of Avila and Saint Augustine. We propose Universal Desired Resources (UDR) as a post-monetary design principle and show that UDR achieves the most radical intersectional intervention ever proposed: by eliminating the material basis of oppression, it simultaneously dissolves all axes of structural inequality identified by intersectional theory (class, race, gender, disability, sexuality, age, nationality). Drawing on Maslow’s hierarchy of needs, we demonstrate that the Liberation Stack satisfies basic human needs universally, liberating every person to pursue self-actualization in real freedom, what we term synthetic liberty: the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy, against the merely negative liberty of neoliberalism. We propose a progressive transition from Universal Basic Income to UDR, leveraging capitalist automation while building commons alternatives. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, and guifi.net confirms that commons-based systems already operate at scale. We conclude that one of the most important engineering challenges of the 21st century is building the decentralized infrastructure through which robotic automation serves all of humanity, enabling a society centered on care, love, and the common good.
@@
Keywords: commons-based peer production; federated systems; cooperative platforms; sociotechnical design; mutual aid; digital sovereignty; open-source AI; mesh networks; platform cooperativism; democratic technology; universal approximation; deep reinforcement learning; full automation; universal desired resources; machine consciousness; computational ethics; intersectionality; Maslow’s hierarchy; synthetic liberty; accelerationist ecologism; nuclear fusion; freedom of faith
@@ 1 Introduction: The Convergence Window
The central claim of this paper is that anarcho-accelerationism is not a partisan ideology but a bridge-building framework. It finds genuine common ground with every major political and intellectual tradition of our time (Section˜3), demonstrates that UDR achieves universal intersectional liberation by eliminating the material basis of all axes of oppression, satisfies the full hierarchy of human needs as described by Maslow, and provides a progressive transition path from the current system through synthetic liberty (Section˜4). It offers a technical architecture (Section˜5) powered by frontier clean energy technologies within an accelerationist ecological framework, grounded in empirical evidence (Section˜6). We believe this matters for computer science because we are the ones building the infrastructure. Every protocol we design, every platform we deploy, every algorithm we optimize embodies a political choice about who has power and who does not (Winner, 1980; Lessig, 2006). We can choose differently.
@@ 2 The Automation Imperative: Why This Framework, Why Now
The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The Liberation Stack (Section˜5) describes existing technologies, but the trajectory points further. Advances in additive manufacturing (3D printing at molecular scales), synthetic biology (DNA-programmable organisms and materials), and recursive AI self-improvement (where AI systems design better AI systems, better robots, better energy systems, and better versions of themselves) converge toward a productive infrastructure capable of synthesizing any physical object, any material, any energy configuration, bounded only by the laws of physics rather than the limits of human labor or capital. 3D bioprinters already fabricate living tissue; DNA origami techniques fold genetic material into nanoscale structures; large language models already write the code that trains the next generation of models. The logical endpoint is not merely automation of existing jobs but the creation of a universal fabrication commons: a decentralized network of robotic systems capable of producing, on demand, whatever any community or individual needs, from food to medicine to housing to energy infrastructure.
@@
Taken together, these four consequences yield a vision that goes well beyond conventional proposals for technology-assisted social reform. We are not proposing a better welfare state, a more humane capitalism, or a fairer distribution of existing wealth. We are proposing that the combination of decentralized full automation, commons governance, and the ontological impossibility of machine suffering makes possible, for the first time in history, a society in which no conscious being is compelled to perform instrumental labor, and every conscious being has access to whatever they need to live a life of care, creativity, and meaning. The remainder of this paper provides the political bridges (Section˜3), the technical architecture (Section˜5), and the empirical evidence (Section˜6) for building this society.
@@ 3 The Bridge-Building Thesis
Crucially, anarcho-accelerationism respects the freedom of faith not merely as a political concession but on principled epistemological and ontological grounds. As argued by Lumbreras and Garrido-Merchán (2024), the ontological scope of theology, the nature of God, the soul, grace, and mystical union, categorically escapes the epistemological reach of artificial general intelligence (AGI). Through the frameworks of Saint Teresa of Avila’s Interior Castle (Las Moradas) and Saint Augustine’s Confessions, Lumbreras and Garrido-Merchán demonstrate that human interiority, the layered structure of the soul oriented toward the divine, is irreducible to computational processes. Saint Teresa maps seven successive moradas (dwelling places) of the soul, each representing a deeper level of self-knowledge and encounter with God, culminating in mystical union. Saint Augustine’s principle that “truth dwells in the inner man” (in interiore homine habitat veritas) grounds genuine knowledge in a first-person subjective perspective that no information-processing system can replicate. AGI can simulate theological discourse, generate text about spiritual experience, and model the behavioral correlates of religious practice, but it cannot undergo the interior journey that Teresa describes, because that journey requires phenomenal consciousness, which, as we have argued (Section˜2), is independent of computational intelligence (Garrido-Merchán and Lumbreras, 2022). The theological domain, precisely because it concerns realities accessible only through interiority and faith, lies beyond the epistemological horizon of any computational system, no matter how powerful. This means that full automation and AGI do not threaten authentic religious life; they liberate time for it. A person freed from instrumental labor by the robotic commons has more time for prayer, contemplation, community worship, charitable service, and the cultivation of the interior life that Teresa and Augustine describe. Anarcho-accelerationism, far from being incompatible with faith, creates the material conditions for the deepest spiritual flourishing.
@@
Faith traditions Common good Shared dignity, mutual aid Operationalization through commons governance; freedom of faith preserved
Intersectionality Liberation from all oppressions Elimination of structural inequality UDR removes material basis of all axes of oppression
@@
4 From Intersectional Liberation to Human Flourishing
The bridge-building thesis (Section˜3) demonstrates that anarcho-accelerationism finds common ground with existing traditions. In this section we go further: we show that decentralized full automation with Universal Desired Resources (UDR) achieves what no previous framework has accomplished, the simultaneous elimination of all axes of structural oppression identified by intersectional theory, the universal satisfaction of basic human needs as described by Maslow’s hierarchy, and the provision of real freedom (not merely formal freedom) through a progressive transition from the current system. This is not incremental reform; it is the material basis for universal human flourishing.
4.1 Intersectional Liberation Through UDR
Intersectionality, as articulated by Crenshaw (1989) and developed by Collins (2019), identifies that systems of oppression do not operate independently but intersect: race, gender, class, sexuality, disability, age, nationality, and other axes of identity produce compound disadvantages that cannot be understood by examining any single dimension in isolation. A Black disabled woman does not experience racism, ableism, and sexism separately; she experiences a unique, compounded form of oppression that is qualitatively different from the sum of its parts. The “matrix of domination” (Collins, 2019) describes how these intersecting systems are sustained by institutional structures, cultural norms, and, crucially, material conditions: access to employment, housing, healthcare, education, food, and social participation.
We argue that UDR, implemented through the decentralized robotic commons, constitutes the most radical intersectional intervention ever proposed, because it eliminates the material basis of all axes of oppression simultaneously. Consider each axis in turn:
  • Class: UDR eliminates class by definition. When every person has direct access to the goods and services they need from the robotic commons, the distinction between those who own the means of production and those who must sell their labor dissolves entirely. There is no proletariat and no bourgeoisie in a society where production is automated and commons-governed.
  • Race and ethnicity: Racial oppression is sustained by material mechanisms: discriminatory hiring, housing segregation, unequal access to education and healthcare, and the wealth gap produced by centuries of extraction. When employment, housing, food, healthcare, and education are provided by the commons independently of market participation, the material leverage through which racial discrimination operates is removed. A person’s race ceases to determine their access to resources.
  • Gender: The gendered division of labor, with its systematic devaluation of care work performed disproportionately by women (Federici, 2012), loses its material foundation when all instrumental labor is automated. Care remains, but it is no longer an economic burden; it is a freely chosen activity in a society organized around care as its central value.
  • Disability: When access to resources is decoupled from productive capacity, the distinction between “able” and “disabled” bodies loses its economic significance. The robotic commons can be designed with universal accessibility as a first-order constraint. Assistive technologies, produced on demand by the commons, become available to all who need them.
  • Sexuality and gender identity: Economic dependence on employers and landlords who may discriminate on the basis of sexual orientation or gender identity is eliminated when material needs are met through the commons. LGBTQ+ individuals no longer face the economic coercion that forces them into closeted lives or precarious employment.
  • Age: Both the young and the elderly, currently marginalized by their exclusion from wage labor, gain full access to the resources of the commons. The elderly are not discarded as “unproductive”; children are not instrumentalized as future workers. Both are valued for what they are: persons deserving of care.
  • Nationality and migration: When material well-being is not tied to citizenship in a wealthy state, the economic desperation that drives forced migration is eliminated. People move freely because they choose to, not because survival requires it.
The key insight is structural: intersectional oppression is maintained through material scarcity administered through markets. When markets are superseded by the robotic commons under UDR, the material infrastructure of oppression is dismantled at its foundation. This does not mean that cultural prejudices disappear automatically; they do not. But it means that prejudice loses its institutional teeth: the capacity to deny resources, opportunities, and dignity on the basis of identity. Cultural transformation follows material transformation, as the historical record of post-emancipation, post-suffrage, and post-civil-rights societies demonstrates, imperfectly but unmistakably.
Table˜4 summarizes how UDR addresses each axis of intersectional oppression.
Table 4: Intersectional liberation through UDR: how the robotic commons eliminates the material basis of each axis of oppression.
Axis of oppression Material mechanism UDR intervention
Class Ownership of means of production Commons ownership eliminates class distinction
Race/ethnicity Discriminatory access to employment, housing, education Universal access independent of market participation
Gender Devaluation of care labor, wage gap All instrumental labor automated; care freely chosen
Disability Exclusion from productive economy Access decoupled from productive capacity
Sexuality Economic coercion via discriminatory employers Material independence from discriminatory institutions
Age Marginalization of non-productive ages Universal access regardless of labor capacity
Nationality Citizenship-based access to resources Material well-being not tied to state membership
4.2 Maslow’s Hierarchy and the Liberation Stack
Abraham Maslow’s hierarchy of needs (Maslow, 1943), despite its well-known limitations as a rigid stage theory, captures a fundamental insight about human motivation: when basic survival needs are unmet, the higher capacities of creativity, intellectual exploration, moral reasoning, and self-actualization remain largely inaccessible. A hungry person does not philosophize; a person in physical danger does not compose music; a person consumed by the struggle for shelter does not pursue scientific inquiry. The hierarchy identifies five levels: physiological (food, water, shelter, warmth), safety (security, stability, health), belonging (love, friendship, community), esteem (recognition, contribution, competence), and self-actualization (creativity, purpose, spiritual growth).
The Liberation Stack (Section˜5) and UDR are designed, whether by explicit intention or structural convergence, to satisfy the first four levels universally, thereby liberating every human being to pursue the fifth. Figure˜4 illustrates the mapping.
Physiological: food, water, shelter, energy Safety: health, stability, ecological security Belonging: community, connection, mutual aid Esteem: contribution, recognition, knowledge Self-actualization: creativity, spirituality, love Layers 0–2: Energy, Manufacturing, Food UDR + Ecological constraints Layers 3, 5: Communication, Governance Layer 4: Knowledge commons Real freedom: the goal
Figure 4: Maslow’s hierarchy of needs mapped onto the Liberation Stack and UDR. The material layers of the stack (energy, manufacturing, food) satisfy physiological needs. UDR within ecological boundaries provides safety. Communication and governance commons enable belonging. Knowledge commons support esteem through contribution. Self-actualization, the pinnacle, is what remains when all other needs are met: creativity, spirituality, love, and the free development of each person’s unique potential.
The material layers of the Liberation Stack (energy, manufacturing, food) ensure that every person’s physiological needs are met: clean energy for warmth and cooking, automated manufacturing for shelter and clothing, agroecological production for nutritious food. UDR, constrained by ecological sustainability, provides safety: the security of knowing that one’s material needs will always be met, regardless of employment status, health, age, or any other contingency. Communication and governance commons provide belonging: the tools for meaningful participation in community life, for maintaining relationships across distance, for democratic self-governance. Knowledge commons provide esteem: the opportunity to contribute, to learn, to be recognized for one’s intellectual and creative work.
What remains, when the first four levels are universally satisfied, is self-actualization: the free development of each person’s unique potential. This is the society that anarcho-accelerationism envisions, not a society of idleness or consumption, but a society where every person is free to pursue the highest expressions of human existence: art, science, philosophy, spirituality, love, care, and the cultivation of community. This is real freedom.
4.3 The Transition Path: From Progressive UBI to UDR
A common objection to transformative visions is that they require an impossible leap from the current system to a radically different one. Anarcho-accelerationism rejects this false dichotomy. The transition from the current system to the care-centered society is progressive, and it leverages the very mechanisms of capitalist development to transcend capitalism itself.
The key insight is that automation is already happening, driven by the profit motive of capitalist firms. Every robot deployed in a factory, every AI system that replaces a white-collar function, every autonomous vehicle that eliminates a driving job, is a step toward full automation. Anarcho-accelerationism does not oppose this process; it redirects it. The fruits of capitalist technological development, the robots, the AI systems, the renewable energy infrastructure, the communication networks, are available for appropriation by the commons. This is not theft; these technologies were built on publicly funded research (Mazzucato, 2013), trained on collectively produced data, and developed using open-source tools created by volunteer communities.
The transition mechanism is a progressive continuum from Universal Basic Income (UBI) to Universal Desired Resources (UDR), illustrated in Figure˜5:
Current system Negative liberty Wage dependence Market access only Progressive UBI Partial protection Cash transfers Market still dominant Expanded UBI + Commons Growing commons Reduced market Synthetic liberty Full UDR Real freedom Post-monetary Commons-governed Automation beginsCommons growFull automationProgressive protection of displaced workers at every stage
Figure 5: The progressive transition from the current system to full UDR. At each stage, workers displaced by automation are protected by progressively expanding mechanisms: first UBI (cash transfers within the market system), then UBI supplemented by growing commons infrastructure, and finally full UDR (direct access to the robotic commons). The transition leverages capitalist technological development while progressively building commons-based alternatives.
Stage 1: Progressive UBI. As automation displaces workers, governments introduce and progressively expand UBI, funded by taxation of automated production, carbon taxes, and wealth taxes. This is not anarcho-accelerationism’s endpoint; it is a transitional mechanism within the existing monetary system. UBI provides a floor of material security but does not challenge the market’s dominance. However, it buys time and creates political space for the construction of commons infrastructure.
Stage 2: UBI + Growing Commons. Simultaneously, communities build the Liberation Stack: community energy cooperatives, fablabs, agroecological food systems, mesh networks, knowledge commons, governance platforms. As commons infrastructure grows, an increasing share of material needs is met outside the market. The effective purchasing power of UBI increases because fewer goods need to be purchased: energy comes from the community microgrid, food from the community farm, communication from the community mesh network. UBI begins to evolve: rather than a monetary transfer, it increasingly functions as a guarantee of access to commons resources.
Stage 3: Full UDR. When the robotic commons can autonomously produce all material goods and services, money becomes unnecessary. UBI is replaced by UDR: every person requests what they need from the productive commons, subject only to ecological sustainability constraints. The transition is complete not through revolution but through the progressive construction of an alternative infrastructure that makes the old system obsolete.
This progressive path embodies what we term synthetic liberty, in contrast to the negative liberty of neoliberalism. Isaiah Berlin’s classic distinction (Berlin, 1969) between negative liberty (freedom from interference) and positive liberty (freedom to act, to realize one’s potential) illuminates the inadequacy of the neoliberal conception of freedom. Neoliberalism offers negative liberty: you are “free” to choose your employer, your health insurance, your housing, but only if you have the money to exercise that choice. A person who is formally free to choose but materially unable to afford any meaningful option is not free in any substantive sense (Sen, 1999). Van Parijs (1995) argues that real freedom requires not just the absence of coercion but the material conditions for genuine self-determination.
Synthetic liberty, as we define it, is the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy. It is “synthetic” in two senses: first, it is constructed (not naturally occurring; it must be built through deliberate institutional and technological design); second, it synthesizes the insights of multiple traditions (liberal autonomy, socialist equality, ecological sustainability, feminist care, communitarian belonging). Synthetic liberty is what remains when the material constraints that make negative liberty a hollow promise are removed by the robotic commons. It is the freedom to create, to care, to learn, to love, to worship, to build community, not because you have “chosen” these activities from a market menu, but because the material conditions of your existence no longer compel you to do otherwise.
@@ 5 The Liberation Stack: A Technical Architecture
Anarcho-accelerationism is not merely a philosophical position. It proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜6. Each layer is composed of existing, mature, open-source technologies.
@@
Full automation of material production requires abundant, cheap, clean energy, and the accelerationist ecological thesis holds that the answer to the energy challenge is not degrowth but directed technological acceleration toward clean abundance (Jacobson et al., 2017). The convergence of multiple frontier energy technologies, all maturing simultaneously in the 2020s, makes this thesis empirically credible for the first time in history. We term this vision accelerationist ecologism: the conviction that ecological sustainability is best achieved not by reducing human activity but by powering it entirely with clean, renewable, and ultimately abundant energy sources that operate within planetary boundaries (Raworth, 2017).
Table˜5 summarizes the frontier energy technologies that underpin the energy layer of the Liberation Stack. These are not speculative proposals; they represent technologies with demonstrated prototypes, substantial investment, and clear pathways to commercial deployment.
Table 5: Frontier energy technologies for accelerationist ecologism: demonstrated innovations from 2020 onward that enable abundant clean energy for full automation.
Technology Key milestone (2020–2026) Projected impact Decentralization potential
Perovskite-silicon tandem solar 34.85% efficiency (LONGi, 2025); printable on flexible substrates (Green et al., 2023) Solar below \$0.01/kWh; panels on every surface Very high: modular, local manufacturing
Enhanced geothermal (EGS) Fervo Energy: first commercial 400 MW plant (2024–2026); 24/7 baseload (Tester et al., 2006) Geothermal anywhere on Earth; 5,000+ GW US potential High: community-scale plants
Nuclear fusion NIF ignition Q=4.13 (2025); CFS SPARC completion (2026); Helion/Microsoft PPA (Creely et al., 2020) Near-unlimited baseload energy by 2035–2040 Medium: regional plants, no carbon
Iron-air batteries Form Energy: 100+ hours storage at <<\$20/kWh; zero thermal runaway (Kwon et al., 2024) Solves multi-day renewable intermittency Very high: uses iron and air
Green hydrogen H2Pro E-TAC: 95%+ efficiency, \$1/kg target (Chatterjee et al., 2021) Replaces fossil fuels in industry, transport High: modular electrolyzers
Long-duration storage DOE \$1.76B for compressed air; gravity storage commercialized (U.S. Department of Energy, 2024) Grid-scale seasonal storage High: geological storage sites
Small modular reactors NuScale NRC-certified; Oklo fast reactor recycles nuclear waste Carbon-free 24/7 bridging technology Medium: factory-built, truck-delivered
[…]
The convergence of these technologies implies a three-phase energy trajectory for the Liberation Stack. In the near term (2025–2030), perovskite tandem solar cells, enhanced geothermal systems, and iron-air batteries provide the technical foundation for community energy sovereignty at dramatically lower costs than today. In the medium term (2030–2040), first-generation commercial fusion plants, green hydrogen at grid parity, and long-duration storage systems enable regional energy abundance sufficient to power full industrial automation. In the long term (2040+), mature fusion technology provides near-unlimited clean baseload energy, while direct air capture powered by this abundance actively reverses historical carbon emissions, healing the atmosphere rather than merely stabilizing it.
Critically, the most decentralizable technologies, solar, enhanced geothermal, iron-air batteries, and green hydrogen, are precisely those that mature earliest. This means that community energy sovereignty does not depend on fusion or other centralized technologies; it can be achieved with technologies that are already commercially available or will be within five years. Fusion and SMRs serve as complementary technologies for energy-intensive applications (heavy industry, materials recycling, large-scale direct air capture) while community-scale renewables provide the distributed foundation. The result is an energy commons that is simultaneously abundant, clean, decentralized, and resilient: precisely what full automation requires, and precisely what the planet needs.
@@
The Liberation Stack is designed for compositional sovereignty: each layer can be adopted independently, but the layers reinforce each other. A community with energy sovereignty has more bargaining power. A community with communication sovereignty cannot be silenced. A community with governance tools can make collective decisions about all other layers. Together, the layers form a complete alternative infrastructure that does not depend on either state provision or market availability. Table˜6 summarizes the architecture.